OpenStudy (deadshot):

How do I find g(f(x)) if f(x) = 2x - 3 and g(x) = x - 1?

OpenStudy (anonymous):

g(DeadShot) = DeadShot - 1 isn't it?

OpenStudy (anonymous):

now what is g(f(x))?

OpenStudy (akashdeepdeb):

g(x) = x-1 So g(f(x) = f(x) - 1 [Now put f(x)'s value in it] Understood? :D

OpenStudy (deadshot):

so then g(f(x)) = (2x - 3) - 1?

OpenStudy (anonymous):

@DeadShot YUP dude ;)

OpenStudy (deadshot):

and simplified it would be 2x - 4, right?

OpenStudy (deadshot):

Thanks!